nandrani:

A water cooler holds 15 liters of sports drink. Approximately how many gallons is this? (1 gallon = 3.785 liters)

nandrani:

I know its 4 gallon but how did they come. up with it.

TheSmartOne:

Hey there! Welcome to QuestionCove!!

TheSmartOne:

The key to solving your question is this little piece of information: 1 gallon = 3.785 liters

TheSmartOne:

This is what's called our conversion factor, basically where we can convert between gallons and liters

nandrani:

15 / 3.785 = 3.9

TheSmartOne:

To use it, you have to be very careful with the units!! To go from one thing to another.. you want to make sure the units of that first thing cancel out and all you're left with it units of the second thing that you're converting to So basically since 1 gallon = 3.785 liters That means (1 gallon)/(3.785 liters) = 1 If you have 5 = 5 Then obviously 5/5 = 1 So when we have 15 liters, we can multiply it by either (1 gallon/3.785 liters) Or multiply it by (3.785 liters)/(1 gallon) Because that fraction equals 1, we're not changing the initial quantity. we're just changing the units, we're just changing what we're measuring with. Instead of liters, we're trying to find how much it is in gallons

TheSmartOne:

Does that make sense so far?

TheSmartOne:

Also, you did it right but it would be 3.963.. And it continues on The question then says **approximately** so you can just round to 4 gallons

TheSmartOne:

\(\sf 15~liters \times \frac{1~gallon}{3.785~liters}\) \(\sf 15~liters \times \frac{3.785~liters}{1~ gallon}\) So which fraction would we want to use? In which one would the liters cancel out and we would be left with gallons in the numerator?

jhonyy9:

@TheSmartOne like an idea try explain it using the rule of 3 simple - probably may be more understandably

jhonyy9:

@nandrani do you understand the above wrote ?

jhonyy9:

by @TheSmartOne

jhonyy9:

1 gal. mean 3,785 l x gal mean 15 l cross multiplie and calcule the x